# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v3

@@ 7.1. Implications for Users
Our analysis reveals that users are not passive recipients of Gen-AI technology but are increasingly engaged, expressive, and collaborative (Observation 3.3). Reviewers provide constructive feedback to co-create better experiences with developers. As Gen-AI becomes more embedded in users’ lives, particularly in the emotional or educational contexts, users come to expect empathy, personalization, and reliability (Implications 1, 3, and 5). This shift highlights the importance of treating users not merely as consumers, but as stakeholders who help shape and refine Gen-AI apps.
Furthermore, our findings highlight the need for further research into the evolving relationship between users and Gen-AI conversational agents. Users increasingly describe these systems as trusted companions or emotional supports. Understanding the psychological and social bonds users form will be crucial for guiding the ethical design, deployment, and regulation of AI in emotionally sensitive contexts (Implication 5).
@@ 7.2. Implications for Platform Owners
Platform owners must bear responsibility for ensuring that Gen-AI apps comply with ethical standards (Observation 3.10). As AI capabilities evolve rapidly, app store operators should implement both technical tools and regulatory frameworks that evaluate whether submitted Gen-AI apps meet baseline criteria for safety, transparency, and responsible data usage. This may include requiring apps to disclose whether AI is used, clarifying how outputs are generated, and verifying that training data sources are ethically obtained. By introducing certification processes or audit mechanisms for Gen-AI features (particularly in apps that claim educational, therapeutic, or professional utility), platforms can promote accountability and mitigate harm, fostering greater trust among users and developers alike.
App platforms play a critical role in shaping the quality of user feedback. Our results show that a large proportion of user reviews are non-informative (Observation 1.2), which limits their value for developers and researchers. Future research could explore how redesigning review interfaces might encourage more meaningful feedback. One possible approach is to prompt users to answer specific questions or rate distinct topics, such as usability, design, or content accuracy, separately from the overall rating.
Beyond improving review quality, our findings also reveal structural limitations in how app stores currently support developer–user interaction. As shown in our developer response analysis (Implication 6), review threads in Gen-AI apps function primarily as an initial point of contact rather than a space for resolving AI-specific failures. Many issues require diagnostic information that cannot be shared publicly or involve model- and policy-level changes that cannot be deployed quickly. Platform owners can play an important role by offering richer, privacy-preserving support channels that complement review sections and allow developers to investigate Gen-AI issues more effectively. Strengthening these communication pathways would improve user experience and reduce the friction developers face when addressing complex AI behaviours.
@@ 7.3. Implications for Developers
Developers face growing pressure to meet user expectations that are increasingly shaped by leading models, such as ChatGPT (Implication 4). Users demand fast, accurate, and context-aware outputs across a wide range of use cases, including education, professional assistance, and emotional support (Implications 3 and 5). Our findings highlight several critical areas for improvement: fine-tuning censorship filters to avoid false positives (Observation 3.9; Implication 2), supporting accessibility and inclusivity concerns (Implication 5), and enabling more personalized and flexible user experiences (Implication 3). Developers should also consider offering modular features and customization options, such as adjustable censorship settings, adjustable AI functionalities, and adaptive content formats. Importantly, developers must think critically about when and how AI is used, ensuring that integration adds meaningful value, as not every problem or task requires an AI solution (Observation 3.7; Implication 7).
@@ 7.4. Implications for Testing Tool Designers
Our LLM-based framework achieves 91% accuracy in topic extraction and assignment (Observation 1.4), demonstrating the effectiveness of prompt design and data pre-filtering (Observations 1.1 and 1.3). These results point to new opportunities in automated review analysis, particularly in testing and quality assurance workflows. Designers of testing tools should consider incorporating LLMs for top topic extraction and assignment. Our methods can help uncover nuanced user feedback across languages, accessibility contexts, and time, significantly enhancing product refinement cycles.
@@ 7.5. Implications for Requirements Engineers
Requirements engineers can benefit from the structured insights our topic analysis framework provides. The ability to track how user expectations evolve by topic over time can inform roadmap planning and feature prioritization (Observation 3.3). Our analysis also reveals that users clearly articulate feature-specific frustrations and preferences, providing actionable guidance for interface design, feature retention, and refinement. Engineers should leverage such review data to validate or revise assumptions about user needs, especially in light of shifting demographics and rising expectations for Gen-AI-powered functionality. Beyond Gen-AI apps, the same framework and analytical tools can be applied to user reviews in other domains, enabling engineers to systematically identify evolving requirements and usability concerns across a wide range of software systems.
@@ 7.6. Implications for Policy Makers and Regulators
The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection. While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples (Observation 3.10), which may indicate normalization rather than resolution. Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced (Implication 2). Furthermore, content moderation must strike a balance between safety and usability. Regulators should develop protocols to mitigate the harm caused by unregulated AI generations (Observation 3.11; Implication 2), particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support.
@@ 8. Threats to Validity
Construct threats to validity refer to the risk that the measures we use in our study do not accurately reflect the concepts we aim to study. First, we use the average app review rating as a proxy for user satisfaction. However, this metric may not fully capture the sentiment expressed in the review text. For example, users sometimes assign high ratings while simultaneously reporting critical issues or limitations. To mitigate this threat, we supplement the quantitative rating data with manual qualitative analysis of review content, ensuring that nuanced sentiments, both positive and negative, are appropriately captured. Second, our methodology relies on LLMs to automatically filter and classify reviews. To control for potential biases introduced by the LLM-based automation, we manually annotate samples and our results demonstrate high agreement, i.e., 91% accuracy, between the LLM-generated classifications and human annotations, supporting the reliability of using LLMs for these tasks. Additionally, we measure inter-rater reliability to confirm the consistency and trustworthiness of the automated classifications.
Internal threats to validity concern the soundness of the procedures used to collect, analyze, and interpret the data. A key internal threat relates to the uncertainty surrounding the timeline for introducing the Gen-AI features app. While we manually verify that all selected apps currently support Gen-AI functionality, we cannot determine the exact date of integration. To reduce the risk of analyzing reviews unrelated to Gen-AI, we exclude all reviews prior to 2022, i.e., the year ChatGPT was released. While this filtering improves temporal alignment, it cannot fully guarantee that all remaining reviews pertain to post-Gen-AI functionality.
External threats to validity concern the ability to generalize the results. Our analysis focuses on publicly available user reviews from the Google Play Store, which may not fully reflect user experiences on other platforms, such as the Apple App Store. This single-platform scope may limit the generalizability of results to other ecosystems with different user bases or review practices. We take several steps to enhance the generalizability of our results. First, we include a diverse set of apps across multiple categories to ensure a wide range of topics and app functionalities. Second, although the selected apps may vary in implementation details from one platform to another, many of the Gen-AI topics discussed (i.e., Content Quality, Content Policy & Censorship, and Accessibility & Inclusivity) are prevalent across platforms and commonly adopted by competing apps within and across categories. Consequently, the insights derived from our analysis of user perceptions and feedback are likely to be transferable to other apps that implement similar Gen-AI capabilities. Developers of comparable apps, including those not represented in our dataset or operating on different platforms, may benefit from these findings. In particular, the results can help them better understand user concerns, inform feature design decisions, and mitigate risks associated with Gen-AI integration.
@@ 9. Conclusion
The rapid integration of Gen-AI features into mobile apps has transformed how users interact with technology. In this study, we present a user-centered analysis of Gen-AI mobile apps by examining over 676,066 reviews from 173 apps on the Google Play Store. Using our proposed four-phase SARA methodology and prompt-based LLM techniques, we uncover actionable insights into how users experience and evaluate Gen-AI functionalities.
Our results show that LLMs can reliably extract user-discussed topics with high accuracy, i.e., 91%, when supported by five-shot prompting and data cleaning. We identify the 10 most frequently discussed topics revealing user behavior and challenges that inform actionable insights for developers and researchers. Furthermore, our findings suggest an evolution in user expectations, with a shift in user behavior, such as shifting from early adopters who were more forgiving of limitations to a mainstream user base that demands higher expectations. Future work will explore cross-platform comparisons and further investigate how expectations vary across different user profiles (e.g., casual vs. expert users) as Gen-AI tools continue and mature.
Data Availability. We provide a replication package, including manually labeled data, and Python scripts to replicate the analyses at omitted.

### v4

@@ 8.1. Implications for Users
Our findings suggest that users are not passive recipients of Gen-AI technology but active participants who shape system behavior through feedback and interaction. This evolving role highlights several key actions for users:
Engage responsibly with AI as both a tool and an emotional companion. Gen-AI systems increasingly support accessibility, wellbeing, and companionship (Opportunity 1), but our findings also show that user relationships can evolve toward emotional dependence and addiction (Finding 3.4). While these systems can provide meaningful support, users should remain aware of the risks of over-reliance and recognize the limitations of AI as a substitute for human relationships. Maintaining balanced usage and critical awareness is essential, particularly in emotionally sensitive contexts.
Develop informed expectations of AI capabilities and limitations. Users should be aware that Gen-AI systems may produce incorrect, inconsistent, or unexpected outputs, particularly for complex tasks. As highlighted in earlier findings, many frustrations stem from mismatched expectations of system capabilities. Understanding these limitations can help users interact more effectively with AI systems and interpret outputs more critically (Challenge 1).
Be mindful of content moderation policies and platform differences. Content moderation practices vary significantly across Gen-AI apps and can affect the type of content users are able to generate. Perceived bias or restrictions may result from underlying safety mechanisms. Being aware of these differences enables users to make informed choices about which platforms align with their expectations regarding bias, privacy, freedom of expression, and content access (Challenge 2 and Finding 3.5).
@@ 8.2. Implications for Platform Owners
Our findings highlight the critical role of platform owners in ensuring the safe deployment of Gen-AI apps, improving feedback quality, and supporting effective developer-user interaction.
Enforce transparency and accountability in Gen-AI systems. Platform owners should ensure that Gen-AI apps comply with ethical and safety standards as capabilities evolve. This includes requiring developers to disclose the use of AI, explain how outputs are generated, and verify that training data sources are ethically obtained (Challenge 2, Finding 3.5). App stores could further support this by providing a certification badge indicating that a Gen-AI app meets AI regulatory standards, including safeguards against over-reliance, addiction, and inappropriate content (Opportunity 1, Finding 3.4). Introducing certification processes or audit mechanisms is important, especially for apps involving education, wellbeing, or young users, as they can improve accountability and reduce potential harm.
Improve the quality and structure of user feedback. Our results show that a large proportion of user reviews are non-informative (Finding 1.2), limiting their usefulness for developers and researchers. Platform owners can address this by redesigning review interfaces to encourage more structured and meaningful input. For example, prompting users to evaluate specific aspects such as usability, content quality, or accuracy, rather than relying solely on free-text reviews, can lead to more actionable feedback.
Enhance support channels for complex Gen-AI issues. Our developer response analysis reveals that review threads often serve only as an initial point of contact rather than a space for resolving AI-specific issues. Many Gen-AI-related problems require additional context (e.g., prompts, outputs, system logs) or involve model- and policy-level changes that cannot be addressed through public replies. Platform owners should provide richer, privacy-preserving support channels that enable more effective communication between users and developers. Strengthening these mechanisms can reduce friction in addressing complex AI behaviors and improve overall user experience.
@@ 8.3. Implications for Developers
Our findings highlight increasing pressure on developers to meet evolving user expectations shaped by leading Gen-AI apps, while ensuring that AI integration remains meaningful, reliable, and aligned with user needs.
Design for collaborative and evolving user needs. Users increasingly engage with Gen-AI systems as collaborative partners that support creativity, ideation, and problem-solving (Opportunity 2). Rather than passively consuming outputs, users iteratively refine prompts and incorporate AI suggestions into their workflows. At the same time, many use cases emerge organically through user experimentation rather than predefined developer intent (Opportunity 3). Developers should therefore monitor how users adapt AI systems in practice and design flexible, extensible systems that support co-creation, iterative interaction, and evolving use cases.
For example, as users increasingly engage with AI, they begin to form emotional bonds; developers must move beyond optimizing conversational quality and towards designing responsible, safe emotional AI systems. This includes incorporating safeguards such as usage awareness features, intervention mechanisms for excessive use, and improved moderation of harmful conversational behaviors, particularly for users relying on AI for emotional support (Finding 3.4).
Ensure purposeful and value-driven integration of Gen-AI. As user expectations are shaped by powerful systems such as ChatGPT, developers face growing pressure to deliver fast, accurate, and context-aware outputs across diverse domains. However, our findings show that not all features benefit from AI integration (Challenge 3). Developers must carefully evaluate when Gen-AI meaningfully enhances the user experience and when simpler solutions are more appropriate. Effective integration requires balancing innovation with reliability, ensuring that AI capabilities provide clear value rather than introducing unnecessary complexity.
Balance moderation and user expectations. Developers must address challenges related to managing user expectations and system limitations (Challenge 1) and balancing content moderation with creative freedom (Challenge 2). Developers should provide clearer feedback on system limitations and offer customization options that support diverse user needs. Content policy should also incorporate greater transparency in data usage and policy communication. Developers should clearly communicate how user data is handled and how moderation decisions are made, helping users better understand system behavior and build trust (Finding 3.5).
@@ 8.4. Implications for Testing Tool Designers
Our LLM-based framework achieves 91% accuracy in topic extraction and assignment (Finding 1.4), demonstrating the effectiveness of prompt design and data filtering. These results point to new opportunities in automated review analysis, particularly in testing and quality assurance workflows. Designers of testing tools should consider incorporating LLMs for top topic extraction and assignment. Our methods can help uncover nuanced user feedback across languages, accessibility contexts, and time, significantly enhancing product refinement cycles (Finding 2.3).
@@ 8.5. Implications for Requirements Engineers
Our findings suggest extending requirements engineering processes to capture emerging quality attributes and design considerations specific to Gen-AI systems. In particular, the identified opportunities highlight the need to explicitly model user-centered and interaction-driven requirements. For example, AI for Accessibility and Wellbeing (Opportunity 1) requires engineers to account for the emotional impact of Gen-AI systems and incorporate safeguards that mitigate risks such as dependency and user vulnerability. Similarly, AI as a Collaborative Creative Tool (Opportunity 2) calls for requirements that support co-creative interaction, treating users as active collaborators rather than passive consumers. The diversity of user-driven applications observed in AI Versatility (Opportunity 3) further emphasizes the need for flexible and adaptive requirements that accommodate evolving and unanticipated use cases.
At the same time, the identified challenges point to critical gaps in current requirements practices. Managing User Expectations and AI Limitations (Challenge 1) highlights the importance of specifying realistic system capabilities and continuously validating them against user expectations. Balancing Content Moderation and Creative Freedom (Challenge 2) introduces the need to formalize requirements around moderation policies, data safety, and privacy while preserving user autonomy. Finally, Strategic Integration of Gen-AI Features (Challenge 3) suggests that requirements engineers must critically evaluate when Gen-AI functionality is appropriate, ensuring that AI is integrated where it provides clear user value rather than as a default solution. While our study focuses on Gen-AI apps, the SARA framework may be adapted to other domains to support requirements analysis using user-generated data.
@@ 8.6. Implications for Policy Makers and Regulators
The emergence of Gen-AI apps raises important questions about ethics, safety, and user protection (Finding 3.5). While some early users voiced concerns about data ethics and copyright in training data, these concerns faded in later samples, which may indicate normalization rather than resolution. Policy makers must ensure that ethical standards around content ownership, training transparency, and accountability are formalized and enforced. Furthermore, content moderation must strike a balance between safety and usability (Challenge 2). Regulators should develop protocols to mitigate the harm caused by unregulated AI generations, particularly as Gen-AI becomes increasingly pervasive in sensitive domains such as education and emotional support (Opportunity 1, and Finding 3.4)
@@ 9. Threats to Validity
Construct threats to validity refer to the risk that the measures we use in our study do not accurately reflect the concepts we aim to study. First, we use the average app review rating as a proxy for user satisfaction. However, this metric may not fully capture the sentiment expressed in the review text. For example, users sometimes assign high ratings while simultaneously reporting critical issues or limitations. To mitigate this threat, we supplement the quantitative rating data with manual qualitative analysis of review content, ensuring that nuanced sentiments, both positive and negative, are appropriately captured. Second, our methodology relies on LLMs to automatically filter and classify reviews. To control for potential biases introduced by the LLM-based automation, we manually annotate samples, and our results demonstrate high agreement, i.e., 91% accuracy, between the LLM-generated classifications and human annotations, supporting the reliability of using LLMs for these tasks. Additionally, we measure inter-rater reliability to confirm the consistency and trustworthiness of the automated classifications.
Internal threats to validity concern the soundness of the procedures used to collect, analyze, and interpret the data. A potential threat arises from uncertainty in identifying the exact timeline of Gen-AI feature integration for each app. To mitigate this, we adopt a conservative, evidence-based approach by manually examining historical release notes and online sources to determine the earliest documented integration date. While this approach may miss undocumented or informally introduced features, it ensures that all included data reflects confirmed Gen-AI functionality. As such, our integration dates represent the earliest verifiable evidence rather than the precise deployment time, prioritizing reliability and reducing the risk of including pre-Gen-AI reviews.
Additionally, a potential threat arises from ambiguity in determining whether an app genuinely incorporates Gen-AI functionality. In particular, some apps reference “AI” without providing generative capabilities, which may lead to misclassification. To mitigate this risk, we adopt a conservative coding strategy in which a mere mention of AI (e.g., “AI Song”) is not considered sufficient evidence of Gen-AI integration unless it is explicitly supported by documented generative functionality (e.g., “AI Song Generator”). This approach prioritizes precision in identifying Gen-AI apps, reducing the likelihood of including apps that do not provide true generative capabilities.
External threats to validity concern the ability to generalize the results. Our dataset is constructed using systematic keyword-based searches on the Google Play Store followed by manual validation. Google Play Store search results are capped, which prevents enumerating all Gen-AI apps for a given keyword. Accordingly, our goal is representativeness rather than exhaustiveness. By querying multiple taxonomy-derived keywords and retaining only apps with verified generative functionality, we obtain a diverse sample covering a wide range of Gen-AI capabilities and usage contexts. While some apps may not be included, our findings are expected to generalize to apps implementing similar Gen-AI functionalities and design choices.
Additionally, user perceptions and behaviors may differ across platforms. To mitigate this, we conduct a cross-platform comparison between Google Play Store and App Store reviews, observing consistent topic patterns across platforms. While this strengthens the generalizability of our findings, differences in platform policies, user demographics, and review mechanisms may still influence the results.
@@ 10. Conclusion
Gen-AI is rapidly transforming mobile apps, yet understanding how users perceive and interact with these systems remains limited. In this paper, we present SARA, a framework for systematically analyzing user reviews of Gen-AI apps, and apply it to over one million reviews across multiple app categories.
Our findings reveal that user discussions of Gen-AI apps surface key opportunities (i.e., AI for Accessibility and Wellbeing, AI as a Collaborative Creative Tool, and AI Versatility) alongside challenges (i.e., Managing User Expectations and AI Limitations, Balancing Content Moderation and Creative Freedom, and Strategic Integration of Gen-AI Features).
Together, these insights inform implications across stakeholders: encouraging users towards more informed and safe interactions; guiding platform owners in enforcing transparency and accountability in Gen-AI systems; helping developers in developing Gen-AI systems for collaborative and evolving user needs; inspiring testing tool designers to incorporate LLMs in automated review analysis; enabling requirements engineers to capture emerging socio-technical requirements; and informing policymakers and regulators in addressing risks and safeguards. We further observe that user engagement evolves over time, from initial novelty-driven interactions to more mature use, where users increasingly identify risks such as dependency and emotional reliance, while expressing reduced concern for issues such as data policy and privacy.
This work contributes both a scalable analysis framework and empirical insights into user perceptions of Gen-AI systems. These findings inform the design of more effective and user-centered Gen-AI apps and highlight the need to incorporate socio-technical considerations into requirements engineering processes. Future work can build on these findings by examining human-AI collaboration in greater depth, particularly focusing on how users interact with Gen-AI systems as co-creative partners and how such interactions shape user expectations, outcomes, and long-term engagement.
Data Availability. We provide a replication package, including manually labeled data, and Python scripts to replicate the analyses at https://github.com/Safwat-UofT/Gen-AI_User_Reviews.
@@ Acknowledgements.
Appendix A Gen-AI App Categories and Keywords
We derive our keyword set from the Gen-AI taxonomy proposed by Gozalo-Brizuela and Garrido-Merchán (Gozalo-Brizuela and Merchan, 2024). We exclude categories that are not relevant to mobile applications (i.e., biotech, brain, drug discovery, and other), and adapt the remaining categories into keyword-based queries for app store search. Table 11 presents the resulting categories and keywords used in our data collection process. Some keywords (e.g.,“Image Editing” and “Business”) are too broad, so we append “Generative AI” for all keywords when searching the Google Play Store.
Category Keywords
Image Image Editing, Artistic Images, Realistic Images
Video Text-to-Video, Video Production, Image-to-Video
3D Text-to-3D
Code and Software Text-to-Code, Text-to-Multilingual Code
Speech Text-to-Speech, AI Understanding, Speech-to-Text
Business Business
Marketing Marketing, Advertisement, Accounting
Gaming Video game Creation
Music Music Generator
Appendix B Prompts
B.1. Filtering Prompt
B.2. Topic Extraction Prompt
Figure 12. Template of the topic extraction prompts for (E1: 0-shot, E2: 3-shot, and E3: 5-shot) used to extract the top topics from a large sample of stage $x$ and app category $y$. “Insert category name” is where we insert the name of the app category $y$. “Insert sample of reviews” is where we list our large sample of reviews, separated by a new line escape.
B.3. Topic Assignment Prompt
Appendix C Topic Categories
Gen-AI Topic Category GRC AGR # Apps # App Categories
AI Performance 158,873 (15%) 4.2 142 9
Utility & Use Cases 118,405 (11%) 4.8 58 4
Content Quality 113,714 (11%) 3.7 142 8
Creative Potential 87,383 (8%) 4.8 96 5
Content Policy & Censorship 22,371 (2%) 2.4 25 4
Features & Functionality 20,200 (2%) 3.6 34 4
Emotional Connection 14,652 (1%) 4.8 9 1
Comparison to Other Apps 1,225 (0.1%) 4.3 17 2
Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
Monetization Methods & Structure 149,255 (14%) 2.2 160 9
Technical Difficulties 126,377 (12%) 2.1 167 9
Other 111,297 (11%) 3.7 162 9
User Interface & Experience 58,885 (6%) 4.3 145 9
Updates & Evolution 35,489 (3%) 2.6 101 8
Customer Support 9,752 (1%) 2.1 116 6
Features & Functionality 2,387 (0.2%) 4.3 12 2